IQ scores are normally distributed with a mean of 100 and a standard deviation of 15. If a certain person has an IQ of 105, what percentage of the population has an IQ lower than hers?
Answer by edjones(7569)

z=(105-100)/15=5/15=0.333
The area of the normal curve below 0.333 = 0.63 or 63% of the population has an IQ lower than hers.
.
Ed